OpenStudy (anonymous):

a lift is ascending with acceleration g/3. what will be the time period of a simple pendulum suspended from the ceiling if its time period in stationary lift is T

OpenStudy (anonymous):

luk in ascending lift g+a=g/3 a=2g/3 rit then a=Aw^2sinwt formula hai now ab time period k lliye mene equation bna li h par resting lift k eaquation smj ni a rhi

OpenStudy (shayaan_mustafa):

i want to give you a quick help. if elevator is ascending then always add g and acceleration of elevator. if elevator is ascending then always subtract g and acceleration of elevator.

OpenStudy (anonymous):

i know this tricks

OpenStudy (anonymous):

bs ap mje resting lift k equation acceleration k terms m btao kaise likhu

OpenStudy (shayaan_mustafa):

simple heena g. time period of the pendulum in rest condition, either in lift or else is T=2pisqrt(L/g)

OpenStudy (shayaan_mustafa):

anything else?

OpenStudy (anonymous):

dusri equation to y bni meri \[g \div3=A ^{2}4\times10\sin 2\pi t \div T\]

OpenStudy (anonymous):

sory not g/3 its 2g/3

OpenStudy (anonymous):

we have to compare if u luk the ques and we should always used the given quantities i think so

OpenStudy (shayaan_mustafa):

thnx JamesJ here. until i am looking someone else.

OpenStudy (shayaan_mustafa):

thnx alot heena.

OpenStudy (jamesj):

If the lift is ascending at a rate of g/3, the mass of the pendulum experiences a net acceleration, a new apparent gravitational force, \( g' \), of \[g' = g + g/3 = \frac{4}{3}g \] You know \[ T = 2\pi \sqrt{L/g} \] Hence what you have to do is find the period \( T ' = 2\pi \sqrt{L/g'} \) in terms of \( T \).

OpenStudy (shayaan_mustafa):

ok thek hy madam.

OpenStudy (anonymous):

lemme solve wait a sec

OpenStudy (anonymous):

but here g'=2g/3 becoz apparent acc= a+g/3

OpenStudy (jamesj):

Think about it. When a lift takes off upwards, i.e., accelerating upwards, you feel heavier, not lighter. Conversely when it is accelerating downwards you feel lighter. Therefore we need to add g/3 to g, not subtract it.

OpenStudy (anonymous):

i agree but not g we will add a we were taught in upwards a=g+a in downwards a=g-a

OpenStudy (anonymous):

no no u r rit i m telling the formula of tension we will add g/3 to initial g and thnx to u too shayaan

OpenStudy (shayaan_mustafa):

oh yes. you are welcome. BTW how did you personaly mesg?

OpenStudy (shayaan_mustafa):

ary kesa bheja dear? mje notification milta hy. me is site pe new hn. mushkil se 1 month hua hy. mje itna kam ni ata.

OpenStudy (anonymous):

okay if you want to through fundamentals then forget everything(formulas n tricks) and think of tension. make a free body diagram of the pendulum Fnet = T - mg Fnet = ma = T - mg m(g/3) = T-mg T=4mg/3 since this tension decides the apparent weight of the bob, and hence the apparent "g".. so apparent g= 4g/3 now,\[T= 2\pi \sqrt{l/g(apparent)}\] therefore,\[T' = \sqrt{3}/2T\]